OpenStudy (anonymous):

I have to have this done by tomorrow the first day of school. I have to show my wont and calculators are not allowed: A police car recieves a radio call to catch a vehicle which is speeding at 80mph down the highway. The police car, which is 12 miles away, drives after it at 108 mph. How long will it take for the police car to catch up?

OpenStudy (anonymous):

distance is rate times time right?

OpenStudy (anonymous):

idk its summer

OpenStudy (anonymous):

the "vehicle" being chases has a 12 mile head start, so its distance will be \(12+80t\) whereas the cop car will have a distance of \(108t\)

OpenStudy (anonymous):

since evidently they meet at the same time, you can finish by setting \[12+80t=108t\] and solve for \(t\) even in the summer, \(D=R\times T\)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

I thought it was Distance = Speed x Acceleration?

OpenStudy (anonymous):

yw

OpenStudy (jdoe0001):

what about in winter? is d = r*t still? =)

OpenStudy (anonymous):

hey, its summer hw, she cant expect us to get them all right, ill just go wuith the first one, since I know how to do it.

OpenStudy (anonymous):

good idea should solve in exactly 2 steps